OpenStudy (anonymous):

What is the solution? 4a+8b=-16 -4a+b=34

OpenStudy (anonymous):

a is -8 and b is 2 i believe

OpenStudy (anonymous):

You need to solve for one variable and then plug it into the opposite equation. Let's start with the 2nd one: -4a+b=34 b=(34+4a) Now you can plug "b" into the opposite equation to solve for just "a." 4a+8(34+4a)=-16 Solve here for "a" and then plug the "a" value you get back into -4a+b=34 to solve for "b."

OpenStudy (anonymous):

.-. thats to much to do

OpenStudy (anonymous):

my way easier plug in the answer i gave it is correct

OpenStudy (anonymous):

.-. to lazy to explain process

OpenStudy (anonymous):

8x34=272 8x4a=32a So... 4a+272+32a=-16 36a+272=-16 36a=-288 a=-8 Plug a = -8 back into -4a+b=34 -4(-8)+b=34 32+b=34 b=2

OpenStudy (anonymous):

to much xD

OpenStudy (anonymous):

When the equations are simple it's easy to guess, but as they get more complex it's good to know the process lol.

OpenStudy (anonymous):

i do something easier

OpenStudy (anonymous):

4a+8b=-16 -4a+b=34 i add them so 4a + -4a cancel 8b+b is 9b ans -16 + 34 is 18 so what is left is 9b=18 so 2=b then you plug it in to get a lol so much easier then what you did

OpenStudy (anonymous):

Agreed, much faster haha

OpenStudy (anonymous):

i'm lazy XD i like short and easy ways